# Evaluating Similitude and Robustness of Deep Image Denoising Models via Adversarial Attack

### v1

@@ Abstract
Deep neural networks (DNNs) have a wide range of applications in the field of image denoising, and they are superior to traditional image denoising. However, DNNs inevitably show vulnerability, which is the weak robustness in the face of adversarial attacks. In this paper, we find some similitudes between existing deep image denoising methods, as they are consistently fooled by adversarial attacks. First, denoising-PGD is proposed which is a denoising model full adversarial method. The current mainstream non-blind denoising models (DnCNN, FFDNet, ECNDNet, BRDNet), blind denoising models (DnCNN-B, Noise2Noise, RDDCNN-B, FAN), and plug-and-play (DPIR, CurvPnP) and unfolding denoising models (DeamNet) applied to grayscale and color images can be attacked by the same set of methods. Second, since the transferability of denoising-PGD is prominent in the image denoising task, we design experiments to explore the characteristic of the latent under the transferability. We correlate transferability with similitude and conclude that the deep image denoising models have high similitude. Third, we investigate the characteristic of the adversarial space and use adversarial training to complement the vulnerability of deep image denoising to adversarial attacks on image denoising. Finally, we constrain this adversarial attack method and propose the L2-denoising-PGD image denoising adversarial attack method that maintains the Gaussian distribution. Moreover, the model-driven image denoising BM3D shows some resistance in the face of adversarial attacks.
@@
Image denoising, Network interpretable, Adversarial attack, Image denoising Adversarial attack, Deep neural networks.
@@ I Introduction
An interesting property of adversarial samples is their good transferability, i.e., adversarial samples designed for one model can successfully fool another model. To the reason for the existence of adversarial samples transferability, there are many opinions but no conclusion has been reached yet. The prevailing claim is that the transferability is created due to the decision boundaries of the samples are correlated between different models[23, 24, 25]. This also suggests that the adversarial samples are not randomly generated, but are traceable subspaces in a high-dimensional space. Some literature[26, 27] attributes transferability to architectural similarity, where the transferability of adversarial samples between two models increases with architectural similarity. Some researches[28, 18] believe that transferability is caused by the high linear correlation between features extracted in different deep neural networks. For the same set of samples, features extracted using one model can be correlated with features extracted using another model by a simple affine transformation.
Existing adversarial attacks[29, 30] and defense techniques[31, 32, 33] uniformly target on classification tasks, while few involving image denoising tasks aim to pave the way for subsequent interference with other classification tasks[34]. Genzel et al.[35] argue that regression tasks, which mapping inputs to high-dimensional signal manifolds, are resilient to adversarial perturbations even under vanilla end-to-end network architectures. Furthermore, deep image denoising is an uninterpretable black-box model, so its stability, robustness, and the affective range of noise are unknown and rarely discussed. The same is true for the model-data hybrid-driven approach, which retains a portion of the network structure that brings uninterpretability, and whose stability, robustness, and the range of noise that can be removed remain to be verified.
@@
(1) We propose image denoising adversarial attack, a novel task for image denoising, with the aim of verifying the robustness of deep image denoising methods. The proposed denoising-PGD adversarial attack method has a significant adversarial effect and exhibits good generalization.
@@
This paper is organized as follows. Section II reviews typical methods of image denoising and the adversarial attack method PGD. Section III describes the denoising-PGD method, its transferability among various denoising models, and the exploration of the adversarial space. The experimental results are presented in Section IV. Section V analyzes and discusses the experimental results and presents some possible conjectures. Finally, Section VI concludes the paper.
@@ II-D PGD adversarial attack
Goodfellow et al.[18] proposed an efficient one-step attack method based on the gradient of deep neural network, called the Fast Gradient Sign Method (FGSM). FGSM uses relationship between the noise map and image pixels as the loss function. The gradient of the loss function is computed by subtracting or adding a fixed value $\varepsilon$ from each pixel depending on its sign. The basic iterative method (BIM)[36] replaces the one-step algorithm with multiple small steps, extending the FGSM to an iterative algorithm. Another iterative version of FDSM is the projection gradient descent(PGD) method[37], which uses projection gradient descent on a negative loss function to generate adversarial samples and can be described as a saddle point optimization problem. Unlike BIM, PGD perturbs the original image randomly before starting the iteration.
@@ III-B Image denoising adversarial attack
Given a noise image $x$, our goal is to generate an adversarial noise image ${x}^{\prime}$ that can fool the deep image denoising model much than applying on a Gaussian noise image of the same noise level and similar distribution. We name this task as image denoising adversarial attack, whose flowchart is showed in Fig. 2.
@@ III-C Image denoising adversarial attack transferability
are collected on the circle centered at the Gaussian noise image and located in the plane whose basis are $n$ and $u$, where $\theta=\left\{\mathrm{k}/\mathrm{N}^{*}2\pi,\mathrm{k}=1,\ldots,\mathrm{N}\right\}<v,n>$ is the inner product of and . Intuitively, we hope to observe the robustness of denoising models in a tiny sphere neighborhood of the Gaussian noise image. But as images are high-dimensional vectors, unified samples may fail to collect the pivot behavior. Thus, we focus on the samples is a specific cross section of the sphere neighborhood, i.e. the circle in the $n,v$ plane. Note that, the direction of $n$ is set as the 0-angle of the circle.
@@
Figure 3: The distribution of the adversarial space and the normal denoising space is shown, and the closer to the edge the stronger the adversarial performance.
Figure 4: Schematic representation of the adversarial region of the denosing-PGD adversarial attack on the DnCNN model for dataset Set12.
In addition, we tested these generated noise samples on the other deep image denoising models, such as the non-blind denoising model ECNDNet, blind denoising models DnCNN-B and RDDCNN with the unfolding model DeamNet and the plug-and-play model DPIR. The adversarial behaviors of these models are observed and recorded in Fig. 5. For all the tested models, the adversarial regions are concentrated in the angular range from 30 degrees to 150 degrees. It means that all the tested models regressed similar denosing functions which leads to similar adversarial regions. It also explains why the denoising-PGD has unexpected excellent transferability among different types of denoising models, as their adversarial spaces are almost overlap with each other. In a deeper observation, the non-blind denoising model (a) and (b) perform similarly, and the blind denoising model (c) and (d) perform similarly. This demonstrates that the adversarial ranges are similar on the same type of deep image denoising networks. The model (e) is the unfolding method and (f) is the plug-and-play method. Although both of them are model-data hybrid-driven methods, their adversarial regions are similar to pure data-driven methods. At a closer look, DPIR has larger adversarial regions.
Figure 5: Example of Set12 generated on the DnCNN model on the rest of the models with adversarial regions demonstrated.
III-D Image denoising adversarial space
@@
Figure 6: A linear combination of two adversarial samples, where the coefficients are traversed to 1 in steps of 0.1.
@@ IV-A The denoising-PGD for image denoising adversarial attack
Figure 7: The denoising-PGD adversarial attack effect graph(house).
@@
Table 1 lists the adversarial effects caused by denoising-PGD on DnCNN for Set12. The adversarial sample of each image behaves significantly worse than the Gaussian noise image during the denoising process in every evaluation metrics. In average, PSNR drops up to 0.807dB, SSIM drops by 0.044, and MAE increases by 0.855.
@@
Further, we use denoising-PGD to generate adversarial samples for a blind denoising model DnCNN-B with the same hyperparameter setting. In Table 2, it shows that the PSNR drop is up to 4.012, the SSIM drop is 0.216, and the MAE difference is 4.082 on DnCNN-B. The adversarial effect is much stronger on DnCNN-B and the visually difference is much obvious with many noise traces remaining on the image. DnCNN-B is less robust comparing to DnCNN.
IV-B The transferability of denoising-PGD
@@
Figure 8: Transferability test of adversarial samples generated by DnCNN network under denoising-PGD(house).Where the test values are PSNR/MAE/SSIM from left to right, respectively.
@@
We show the denoising performance of the adversarial samples generated on the DnCNN model using denoising-PGD on the Set12 dataset at noise level 25 on the non-blind denoising method FFDNet, the blind denoising method DnCNN-B, and the plug-and-play method DPIR in Table 3. For the averaged PSNR, FFDNet decreases by 0.7579 dB, DnCNN-B decreases by 0.6653 dB, and DPIR decreases by 0.9190 dB. The SSIM and MAE metrics show similar trends to each other. It can be seen that the blind denoising model is relative resistant to the adversarial perturbation created for the non-blind denoising model. However, the plug-and-play model, as a model-data hybrid-driven method, is unexpectedly sensitive to the adversarial perturbation for the non-blind denoising model. Note that, denoising-PGD, which is supposed to have poor transferability, exhibits excellent transferability among multiple deep image denoising methods which are designed in various motivations, mathematical backgrounds, and across years. It indicates that current deep denoising models are essentially similar, at least similar in the local function regression around all the test data.
@@
In the experiments to test the transferability from DnCNN-B, we selected the non-blind denoising methods FFDNet and ECNDNet, the blind denoising methods DnCNN-B and RDDCNN-B, the plug-and-play method DPIR and the deep unfolding method DeamNet to test the adversarial samples generated by DnCNN-B under denoising-PGD. The results are compiled in Table 4.
@@
Comparing to the non-blind denoising method and plug-and-play method with less than 1 dB decrease in PSNR, the blind denoising method DnCNN-B decreases by 4.0114 dB and RDDCNN-B decreases by 2.6097 dB. The SSIM and MAE metrics also show the same trend, i.e., the adversarial attack samples generated with the blind denoising model DnCNN-B are highly transferable on the blind denoising model and is also transferable on non-blind denoising methods as well as plug-and-play methods, but the attack effect is greatly weakened. A possible reason is that blind denoising methods have larger adversarial space comparing to non-blind denoising and plug-and-play methods. Thus, adversarial samples generated for DnCNN-B can fall out of the adversarial regions of non-blind denoising and plug-and-play methods.
IV-C Color image denoising adversarial attack

### v2

@@ Abstract
Deep neural networks (DNNs) have shown superior performance comparing to traditional image denoising algorithms. However, DNNs are inevitably vulnerable while facing adversarial attacks. In this paper, we propose an adversarial attack method named denoising-PGD which can successfully attack all the current deep denoising models while keep the noise distribution almost unchanged. We surprisingly find that the current mainstream non-blind denoising models (DnCNN, FFDNet, ECNDNet, BRDNet), blind denoising models (DnCNN-B, Noise2Noise, RDDCNN-B, FAN), plug-and-play (DPIR, CurvPnP) and unfolding denoising models (DeamNet) almost share the same adversarial sample set on both grayscale and color images, respectively. Shared adversarial sample set indicates that all these models are similar in term of local behaviors at the neighborhood of all the test samples. Thus, we further propose an indicator to measure the local similarity of models, called robustness similitude. Non-blind denoising models are found to have high robustness similitude across each other, while hybrid-driven models are also found to have high robustness similitude with pure data-driven non-blind denoising models. According to our robustness assessment, data-driven non-blind denoising models are the most robust. We use adversarial training to complement the vulnerability to adversarial attacks. Moreover, the model-driven image denoising BM3D shows resistance on adversarial attacks.
@@
image denoising, adversarial attack, denoisng-PGD, robustness similitude.
@@ I Introduction
An interesting property of adversarial samples is their transferability, i.e., adversarial samples designed for one model can successfully fool another model. To the reason for the existence of adversarial samples transferability, there are many opinions but no conclusion has been reached yet. The prevailing claim is that the transferability is created due to the decision boundaries of the samples are correlated between different models[23, 24, 25]. This also suggests that the adversarial samples are not randomly generated, but are traceable subspaces in a high-dimensional space. Some literature[26, 27] attributes transferability to architectural similarity, where the transferability of adversarial samples between two models increases with architectural similarity. Some researches[28, 18] believe that transferability is caused by the high linear correlation between features extracted in different deep neural networks. For the same set of samples, features extracted using one model can be correlated with features extracted using another model by a simple affine transformation.
Existing adversarial attacks[29, 30] and defense techniques[31, 32, 33] mostly target on classification tasks, while few involving image denoising tasks aim to pave the way for subsequent interference with other classification tasks[34]. Genzel et al.[35] argue that regression tasks, which mapping inputs to high-dimensional signal manifolds, are resilient to adversarial perturbations even under vanilla end-to-end network architectures. Furthermore, deep image denoising is an uninterpretable black-box model, so its stability, robustness, and the affective range of noise are unknown and rarely discussed. The same is true for the model-data hybrid-driven approach, which retains a portion of the network structure that brings uninterpretability, and whose stability, robustness, and the range of noise that can be removed remain to be verified.
@@
(1) We propose image denoising adversarial attack, a task for image denoising, with the aim of verifying the robustness of deep image denoising methods. The proposed denoising-PGD adversarial attack method has a significant adversarial effect and exhibits good generalization.
@@
This paper is organized as follows. Section II reviews typical methods of image denoising and the adversarial attack method PGD. Section III describes the denoising-PGD method, its transferability among various denoising models, the exploration of the adversarial space, and the definition of robustness similitude. The experimental results are presented in Section IV. Section V analyzes and discusses the experimental results and presents some possible conjectures. Finally, Section VI concludes the paper.
@@ II-D PGD adversarial attack
Goodfellow et al.[18] proposed an efficient one-step attack method based on the gradient of deep neural network, called the Fast Gradient Sign Method (FGSM). FGSM uses relationship between the noise map and image pixels as the loss function. The gradient of the loss function is computed by subtracting or adding a fixed value $\varepsilon$ from each pixel depending on its sign. The basic iterative method (BIM)[36] replaces the one-step algorithm with multiple small steps, extending the FGSM to an iterative algorithm. Another iterative version of FGSM is the projection gradient descent (PGD) method[37], which uses projection gradient descent on a negative loss function to generate adversarial samples and can be described as a saddle point optimization problem. Unlike BIM, PGD perturbs the original image randomly before starting the iteration.
@@
Recently, Yan et al.[39] proposed OBSATK, a two-step adversarial attack method for evaluating the robustness of deep denoising models, aiming to obtain a bounded perturbation that can decrease the quality of the denoised image. In the experiments, DnCNN-B were demonstrated to be vulnerable to adversarial attacks. Different from this work, we aim to explore characteristics of deep denoising models and the correlation between models from the adversarial attack perspective, rather than finding stronger attack methods.
@@ III-B Image denoising adversarial attack
Given a noise image $x$, our goal is to generate an adversarial noise image ${x}^{\prime}$ that can fool the deep image denoising model to result much lower denoising quality than functioning on a Gaussian noise image with the same noise level and similar distribution. We name this task as image denoising adversarial attack, whose flowchart is showed in Fig. 2.
@@ III-C Image denoising adversarial attack transferability
[…]
@@
Figure 3: Radar diagram of the distribution of the adversarial space and normal denoising space. The closer green area to the edge the stronger is the adversarial performance.
Figure 4: Adversarial regions of the DnCNN model under denoising-PGD attack on the dataset Set12.
In addition, we tested these generated noise samples on the other deep image denoising models, such as the non-blind denoising model ECNDNet, blind denoising models DnCNN-B and RDDCNN-B, the unfolding model DeamNet, and the plug-and-play model DPIR. The adversarial behaviors of these models are observed and recorded in Fig. 5. For all the tested models, the adversarial regions are concentrated in the angular range from 30 degrees to 150 degrees. It means that all the tested models regressed similar denoising functions which leads to similar adversarial regions. It also explains why the denoising-PGD has unexpected excellent transferability among different types of denoising models, as their adversarial spaces are almost overlap with each other. In a deeper observation, the non-blind denoising model (a) and (b) perform similarly, and the blind denoising model (c) and (d) perform similarly. This demonstrates that the adversarial ranges are similar on the same type of deep image denoising networks. The model (e) is the unfolding method and (f) is the plug-and-play method. Although both of them are model-data hybrid-driven methods, their adversarial regions are similar to pure data-driven methods. At a closer look, DPIR has larger adversarial regions.
Figure 5: Adversarial regions of various models under adversarial samples generated for DnCNN.
III-D Denoising model robustness and robustness similitude
We evaluate the denoising model robustness (DMR) as
$DMR=\frac{1}{M}\sum_{i=1}^{M}E(D(x_{i}^{\prime}))-E(D(x_{i}))$ (8)
where $X=\{x_{1},\ldots,x_{M}\}$ is the evaluation dataset and $x_{i}^{\prime}=adv(x_{i},D)$ is the adversarial sample. DMR calculates the average metric change caused by applying adversarial attack to the target model on a specified dataset.
As demonstrated in the previous subsection, all the tested models regress a similar denoising function or at least similar in the local function regression around all the test data. To evaluate the similarity across models, or precisely the local similarity around all the test data, we propose an indicator called denoising model robustness similitude (DMRS):
$DMRS=\frac{1}{M}\sum_{i=1}^{M}\text{IoU}(\text{Radar}(D_{1}(s_{i})),\text{Radar}(D_{2}(s_{i})))$ (9)
where IoU is the Intersection over Union, Radar is the radar diagram calculated on the $n,v$-plane cross section of the sample neighborhood, i.e.
$\text{Radar}(D(s_{i}))=\text{enclosed area of }\left\{\frac{E(D(s_{ij}))}{\max_{j}E(D(s_{ij}))}\right\}$ (10)
where $s_{ij}$ is the noise sample of $x_{i}$ at $\theta_{j}=j/N*2\pi$. Intuitively, DMRS is the averaged IoU between two different models $D_{1}$ and $D_{2}$ over the evaluation dataset. High DMRS indicates two models are essentially similar observed from the robustness perspective.
III-E Image denoising adversarial space
@@
Figure 6: A linear combination of two arbitrary adversarial samples, where the coefficients are traversed to 1 in steps of 0.1.
@@ IV-A The denoising-PGD for image denoising adversarial attack
Figure 7: The denoising-PGD adversarial attack effect (house).
@@
Table I lists the adversarial effects caused by denoising-PGD on DnCNN for Set12. The adversarial sample of each image behaves significantly worse than the Gaussian noise image during the denoising process in every evaluation metrics. In average, PSNR drops up to 0.8 dB, SSIM drops by 0.04, and MAE increases by 0.8.
TABLE I: Comparison of the denoising effect on Gaussian noise images and adversarial noise images for DnCNN model.
Gauss
Adv
Gauss
Adv
Gauss
Adv
Image
PSNR
PSNR(RES)
MAE
MAE(RES)
SSIM
SSIM(RES)
Cameraman
30.0
29.1(-0.9)
5.1
6.3(+1.2)
0.87
0.81(-0.06)
House
33.0
31.4(-1.6)
4.0
5.1(+1.1)
0.86
0.79(-0.07)
Pepper
30.7
29.6(-1.1)
5.3
6.2(+0.9)
0.88
0.82(-0.06)
Fishstar
29.4
28.7(-0.7)
6.3
7.0(+0.7)
0.87
0.85(-0.02)
Monarch
30.3
29.6(-0.7)
5.3
6.0(+0.7)
0.92
0.87(-0.05)
Airplane
29.1
28.6(-0.5)
5.8
6.4(+0.6)
0.87
0.84(-0.03)
Parrot
29.5
28.7(-0.8)
5.8
6.7(+0.9)
0.86
0.81(-0.05)
Lena
30.9
29.9(-1.0)
5.0
6.0(+1.0)
0.88
0.82(-0.06)
Barbara
28.5
28.1(-0.4)
6.9
7.3(+0.4)
0.83
0.82(-0.01)
Ship
29.0
28.3(-0.7)
6.4
7.2(+0.8)
0.83
0.79(-0.04)
Man
28.8
28.2(-0.6)
6.7
7.4(+0.7)
0.81
0.78(-0.03)
Couple
28.7
27.9(-0.8)
6.8
7.7(+0.9)
0.83
0.79(-0.04)
Avg.
29.8
29.0(-0.8)
5.8
6.6(+0.8)
0.86
0.82(-0.04)
@@
Further, we use denoising-PGD to generate adversarial samples for a blind denoising model DnCNN-B with the same hyperparameter setting. In Table II, it shows that the PSNR drop is up to 4.0 dB, the SSIM drop is 0.22, and the MAE difference is 4.1 on DnCNN-B. The adversarial effect is much stronger on DnCNN-B and the visually difference is much obvious with many noise traces remaining on the image. DnCNN-B is less robust comparing to DnCNN.
TABLE II: Comparison of the denoising effect on Gaussian noise images and adversarial noise images for DnCNN-B model.
Gauss
Adv
Gauss
Adv
Gauss
Adv
Image
PSNR
PSNR(RES)
MAE
MAE(RES)
SSIM
SSIM(RES)
Cameraman
30.0
25.1(-4.9)
5.1
10.4(+5.3)
0.87
0.56(-0.31)
House
32.9
27.1(-5.8)
4.0
8.4(+4.4)
0.86
0.58(-0.28)
Pepper
30.7
26.1(-4.6)
5.3
9.5(+4.2)
0.88
0.63(-0.25)
Fishstar
29.3
25.4(-3.9)
6.3
10.3(+4.0)
0.87
0.70(-0.17)
Monarch
30.4
27.2(-3.2)
5.2
8.2(+3.0)
0.92
0.74(-0.18)
Airplane
29.1
25.4(-3.7)
5.8
10.3(+4.5)
0.87
0.61(-0.26)
Parrot
29.4
24.7(-4.7)
5.8
11.1(+5.3)
0.86
0.57(-0.29)
Lena
30.8
26.3(-4.5)
5.1
9.2(+4.1)
0.88
0.64(-0.24)
Barbara
28.3
25.8(-2.5)
7.0
9.7(+2.7)
0.83
0.71(-0.12)
Ship
29.0
25.6(-3.4)
6.4
10.4(+4.0)
0.83
0.64(-0.19)
Man
28.8
28.2(-0.6)
6.7
10.2(+3.5)
0.81
0.65(-0.16)
Couple
28.7
25.1(-3.6)
6.8
10.8(+4.0)
0.83
0.65(-0.18)
Avg.
29.8
25.8(-4.0)
5.8
9.9(+4.1)
0.86
0.64(-0.22)
IV-B Denoising model robustness evaluation
We use denoising-PGD to generate adversarial samples against the target model on Set12. PSNR is used as evaluation metric, as shown in Table III. We can see that data-driven non-blind denoising models are the most robust, followed by hybrid-driven models, while blind denoising models are the least robust.
TABLE III: Denoising model robustness on Set12.
Gauss
Adv
Gauss
Adv
Gauss
Adv
Image
DMR
PSNR
PSNR
MAE
MAE
SSIM
SSIM
DnCNN
-0.94
30.40
29.47
5.39
6.28
0.862
0.819
ECNDNet
-0.87
30.38
29.51
5.40
6.25
0.861
0.814
DnCNN-B
-4.53
30.31
25.78
5.44
9.91
0.858
0.610
RDDCNN-B
-3.80
30.39
26.50
5.46
8.96
0.859
0.665
DPIR
-2.14
30.94
28.80
4.99
7.05
0.873
0.793
DeamNet
-1.27
30.80
29.53
5.12
6.32
0.871
0.799
IV-C The transferability of denoising-PGD
@@
Figure 8: Transferability of adversarial samples generated by DnCNN to various models (house). The test values are PSNR/MAE/SSIM from left to right.
@@
We show the denoising performance of the adversarial samples generated on the DnCNN model using denoising-PGD on the Set12 dataset at noise level 25 on the non-blind denoising method FFDNet, the blind denoising method DnCNN-B, and the plug-and-play method DPIR in Table IV. For the averaged PSNR, FFDNet decreases by 0.7 dB, DnCNN-B decreases by 0.7 dB, and DPIR decreases by 0.9 dB. The SSIM and MAE metrics show similar trends to each other. It can be seen that the blind denoising model is relative resistant to the adversarial perturbation created for the non-blind denoising model. However, the plug-and-play model, as a model-data hybrid-driven method, is unexpectedly sensitive to the adversarial perturbation for the non-blind denoising model. Note that, denoising-PGD, which is supposed to have poor transferability, exhibits excellent transferability among multiple deep image denoising methods which are designed in various motivations, mathematical backgrounds, and across years. It indicates that current deep denoising models are essentially similar, at least similar in the local function regression around all the test data.
TABLE IV: Transferability of adversarial samples generated by DnCNN to FFDNet, DnCNN-B, and DPIR on Set12.
Image Cameraman House Pepper Fishstar Monarch Airplane Parrot Lena Barbara Ship Man Couple Avg.(RES)
FFDNet
PSNR 29.8 33.5 30.9 29.2 30.1 29.0 29.1 31.0 28.7 29.0 28.8 28.8 29.8
Gauss SSIM 0.87 0.87 0.88 0.86 0.92 0.87 0.85 0.89 0.84 0.83 0.81 0.84 0.86
MAE 5.4 3.7 5.1 6.4 5.3 5.9 6.1 4.9 6.6 6.5 6.7 6.7 5.8
PSNR 29.2 31.8 29.9 28.7 29.4 28.4 28.8 30.1 28.3 28.2 28.2 28.1 29.1(-0.7)
Adv SSIM 0.80 0.79 0.82 0.85 0.87 0.83 0.81 0.84 0.83 0.78 0.78 0.80 0.82(-0.4)
MAE 6.2 4.9 6.0 6.9 6.2 6.5 6.6 5.7 7.2 7.4 7.4 7.5 6.6(+0.8)
DnCNN-B
PSNR 29.1 33.2 30.4 29.1 30.2 28.9 28.5 30.9 28.4 28.8 28.8 28.6 29.6
Gauss SSIM 0.83 0.86 0.87 0.86 0.91 0.86 0.83 0.88 0.83 0.82 0.81 0.83 0.85
MAE 5.9 3.9 5.5 6.5 5.3 5.9 6.6 5.1 6.9 6.6 6.8 6.9 6.0
PSNR 28.0 32.3 29.5 28.2 29.9 28.3 27.3 30.4 28.1 28.3 28.4 28.1 28.9(-0.7)
Adv SSIM 0.76 0.83 0.82 0.83 0.89 0.81 0.75 0.86 0.81 0.80 0.79 0.81 0.81(-0.4)
MAE 6.8 4.5 6.2 7.3 5.7 6.7 7.8 5.4 7.3 7.1 7.2 7.4 6.6(+0.6)
DPIR
PSNR 30.1 34.1 31.2 29.8 30.7 29.4 29.3 32.4 29.9 29.2 29.0 29.1 30.3
Gauss SSIM 0.87 0.87 0.89 0.88 0.93 0.87 0.86 0.90 0.88 0.84 0.82 0.85 0.87
MAE 5.3 3.4 5.0 6.0 4.9 5.5 6.0 4.6 5.8 6.2 6.5 6.4 5.5
PSNR 29.3 31.8 29.9 29.3 29.7 28.7 28.8 30.2 29.2 28.5 28.4 28.4 29.4(-0.9)
Adv SSIM 0.78 0.78 0.81 0.86 0.86 0.83 0.80 0.83 0.85 0.79 0.78 0.81 0.81(-0.6)
MAE 6.3 5.0 6.1 6.4 6.0 6.2 6.7 5.7 6.5 7.1 7.2 7.2 6.4(+0.9)
@@
In the experiments to test the transferability from DnCNN-B, we selected the non-blind denoising methods FFDNet and ECNDNet, the blind denoising methods DnCNN-B and RDDCNN-B, the plug-and-play method DPIR and the deep unfolding method DeamNet to test the adversarial samples generated by DnCNN-B under denoising-PGD. The results are compiled in Table V.
@@
Comparing to the non-blind denoising method and plug-and-play method with less than 1 dB decrease in PSNR, the blind denoising method DnCNN-B decreases by 4.01 dB and RDDCNN-B decreases by 2.61 dB. The SSIM and MAE metrics also show the same trend, i.e., the adversarial attack samples generated with the blind denoising model DnCNN-B are highly transferable on the blind denoising model and is also transferable on non-blind denoising methods as well as plug-and-play methods, but the attack effect is largely weakened. A possible reason is that blind denoising methods have larger adversarial space comparing to non-blind denoising and plug-and-play methods. Thus, adversarial samples generated for DnCNN-B can fall out of the adversarial regions of non-blind denoising and plug-and-play methods.
TABLE V: Transferability of adversarial samples generated by DnCNN-B to various models on Set12
Gauss
Adv
Gauss
Adv
Gauss
Adv
Image
PSNR
PSNR
MAE
MAE
SSIM
SSIM
FFDNet
29.66
29.73
5.88
5.89
0.858
0.858
ECNDNet
29.74
29.63
5.86
5.98
0.858
0.854
DnCNN-B
29.78
25.77
5.80
9.88
0.857
0.641
RDDCNN-B
29.60
26.99
5.96
8.40
0.852
0.717
DPIR
30.25
30.15
5.46
5.58
0.871
0.866
DeamNet
30.11
29.96
5.60
5.76
0.868
0.861
IV-D Denoisng model robustness similitude evaluation
We use denoising-PGD to generate adversarial samples on the Set12 dataset for the DnCNN model and then calculate corresponding noise samples at different angles as the evaluation noise sample set for the similitude evaluation of all the models. As shown in Table VI, deep image denoising models of the same type have up to 90% similarity even though the comparison models were proposed across many years. And surprisingly, the model-data hybrid-driven models with partial interpretability also has around 90% similarity with the non-blind denoising models.
TABLE VI: Denoising model robustness similitude on Set12 (DnCNN adv. sample based).
DMRS
DnCNN
ECNDNet
DnCNN-B
RDDCNN-B
DeamNet
DPIR
DnCNN
1
0.97
0.51
0.56
0.93
0.88
ECNDNet
0.97
1
0.51
0.56
0.94
0.90
DnCNN-B
0.51
0.51
1
0.83
0.52
0.54
RDDCNN-B
0.56
0.56
0.83
1
0.58
0.60
DeamNet
0.93
0.94
0.52
0.58
1
0.92
DPIR
0.89
0.90
0.54
0.60
0.92
1
IV-E Color image denoising adversarial attack